# Improving Code Reviewer Recommendation: Accuracy, Latency, Workload, and Bystanders

### v1

@@ 4.1. Design for RevRecV2
Prior works have mostly ranked reviewers based on a set of features and rules, e.g., (Asthana et al., 2019; Mirsaeedi and Rigby, 2020). Since reviewer recommendation is a ranking problem, we select learn-to-rank approach (Burges, 2010), LambdaMART. A pairwise comparison of each reviewer is made and this ordering is then used to determine the final order. In our case we use MART (Multiple Additive Regression Trees) as function to determine the ordering of reviewers. At Meta there is infrastructural support for LambdaMART models at scale using measures collecting data from internal databases on code review and coding.
To train RevRecV2 we used the features shown in Table 2. We trained the model on over 4.3 million instances of these features for reviewers. We evaluated the model on 1.1 million recommended candidates. The feature importance can be seen in Table 2. We discuss A/B test experiment outcomes in Section 5.1 reporting both accuracy and latency results.
@@ 4.2. Design for RevRecWL
At Microsoft (Asthana et al., 2019), they found that initial feedback from developers indicated that reviewers from other projects complained that they were being incorrectly recommended. We made three compromises to convince the engineering team to run the experiment: (1) we set $0.1$ which reduces the influence of workload favoring expertise. (2) we only re-rank the top 5 suggestions from RevRecV2. (3) instead of assigning the top ranked reviewer, we allow the author to select one of the top5 reviewers. These have two important guaranties, first, people with low workload, but little experience will not be inadvertently added to the list eliminating the issue that Microsoft experienced. Second, the Top5 accuracy of RevRecWL = RevRecV2. Since the top 5 developers are displayed Phabricator, this means in the most extreme case, the top developer from RevRecV2 will be displayed as the 5th developer if they have high workload, but the same reviewers will be displayed only the order will change.
@@
(2) $\text{{RevRecWL}}(r)=\frac{\text{{RevRecV2}}(r)}{\text{Workload}(r)}$
Instead of choosing a single measure of workload, we experiment with three types of workload weights, two are novel: The number of upcoming meetings in the next 7 days, the reviewer activity in the last 7 days (Asthana et al., 2019), and time spent reviewing in the last 7 days. We calculate each over a 7 day period because developer patterns on a particular day tend to vary, but the entire work week tends to be more predictable (Chen et al., 2022). In Section 6, we present our historical analysis to determine the best workload measure and our A/B test to determine if we can use RevRecWL in production.
@@ 4.3. Design for RecBystander
The bystander effect is a well known psychological problem that when a group of people are present, it is less likely that any individual will take action (Fischer et al., 2011). At Meta, developers work in a monorepo, so there are times when the author does not know which engineer to assign the review to. In other cases, the author might not mind and anyone in on a particular team can do the review. We decided to use the recommenders we developed in the previous two sections to automatically assign a reviewer.
We experimented with two recommenders in production. BystanderRecRnd randomly assigns one of the top 3 reviewers from RevRecV2 to the review. BystanderRecWL builds on RevRecWL and takes the top reviewer with the a low workload and assigns them to the review. In both cases, we expected to see a decrease in the time-in-review. The results are presented in Section 7.
@@ 4.4. Goal and Guardrail Outcomes Metrics
The vast majority of reviewer recommendation work does not release the recommender in production and only evaluate the accuracy on on historical data. In contrast, the culture at Meta uses a sophisticated A/B test of every new feature and change. The team defines goal and guardrail metrics in advance of the experiment and then a double-blind randomized trail is used to test the existing feature, control, with the new feature, test. Goal metrics are the metrics we hypothesize will improve with the new feature. In contrast, guardrail metrics are the outcome metrics we want do not want to negatively impact. In Table 1, we introduce the the metrics that we use as outcomes in this paper: TopN accuracy, time-in-review, TimeSpent, Latency, Clicks on recommendations, and Workload. Depending on the experiment these outcomes will either be goal metrics or guardrails. For example, in the first experiment with RevRecV2 our goal is to improve TopN accuracy and reduce latency. Our guard metrics are to ensure that reviews do not rubberstamp reviews and rush them, i.e. TimeSpent does not decrease. We discuss the our goals and guardrails metrics in the experimentation sections below.
5. RevRecV2 in Production
@@
In Section 3, we examined the code reviewer recommender literature. We use features from each category of measure we identified in the literature: code ownership, reviewer experience, reviewer participation rate, and familiarity with the author. While prior work have used how often a candidate reviewer in the past, we use a novel measure of how long an reviewer has viewed the author’s changes in the last 90 days TimeSpent.
@@
We conducted an A/B test experiment to understand if the new model is effective in production. Our experimental unit was per diff review. Diffs in the control group, A, continued to receive reviewer recommendations from RevRecV1, the existing model in production. Diffs in the test group, B, received reviewer recommendations from our new model. Our experiment ran between in the late Spring of 2022, on 82k diff which were evenly and randomly split between the two conditions.
@@
Our goal metrics, the metrics we wanted to improve, where the top 1, top 2 and top 3 recommendation accuracy. TopN accuracy in this context is the fraction of times where at least one person belonging to the top n predictions has either commented or acted, e.g., accepted, the diff. We also wanted to improve the latency on the slowest diffs and tracked the latency at the 90th and 99th percentiles.
@@
Our guardrail metrics, the metrics we did not want to negatively influence, were the time diffs were in review (cycle time) and the TimeSpent that reviewers looked at each diff. We did not want to increase cycle time and we did not want recommended engineers to rubberstamp diffs and spend less time looking at them.
5.1. Experimental results for RevRecV2
In Table 3, we see that the new model is much more accurate than RevRecV1. Looking at top-1 accuracy, top-2 accuracy and top-3 accuracy we see 9.14, 12.53 and 14.19 percentage point improvements, respectively, over RevRecV1. In Table 2, we show the feature importance. The most important features that contribute to this accuracy improvement relate to the familiarity of the author with the review (TimeSpent on the author’s prior diffs, 33.89%) and the reviewer participation rate (number of diffs assigned 31.28%). In contrast, code and review ownership are still useful but much less important. The number of files in the current diff that the candidate reviewer has authored in the past accounts for only 4.91%. The number of diffs that a reviewer has acted on and rejected have an importance of 10.10% and 11.56%. Our findings clearly show that being a more active reviewer and being often reviewing the authors diffs are much stronger predictors of who will actually perform a review than code ownership.
Table 3. Improvement in Accuracy for RevRecV2
Top-1 Top-2 Top-3
RevRecV1 43.40 54.36 59.05
RevRecV2 52.54 66.89 73.24
Improvement 9.14 points 12.53 points 14.19 points
When comparing latency of the new model to RevRecV1, we see major wins in Table 4. The 90th percentile latency decreased from 4.43s to 0.3s (a 14x improvement), and p99 latency decreased from 12.37s to 1.26s (a 9x improvement). This reduction in latency largely because our new model uses measures that are stored in a database and does not require blame information to be calculated on the fly.
@@ 5.1. Experimental results for RevRecV2
We also see that diffs in the test condition have more usage/clicks of the suggested reviewer feature, with a 27.3% increase in the number of diffs where a suggested employee reviewer was added using the recommender reviewer component on the diff review page, see Table 5. As well as a 9% increase in in the number of recommendations made. This is partially a result of higher accuracy predictions leading to more clicks as well as lower latency allowing authors to see the recommendations sooner.
Based on these results, RevRecV2 has been rolled out to 100% of engineers at Meta since early Summer 2022.
Table 5. We report the number of diffs that the models were able to create recommendations for and how often the author felt that they were useful, i.e. clicked the recommended reviewer.
@@
In the A/B test, we saw a top 3 accuracy of 73% a 14 point improvement over the previous model. We also saw a 14 times reduction in latency at p90. The recommendations were selected by authors 27% more often with the new model. We did not see any statistically significant changes in the review cycle time or the reviewer viewing time. RevRecV2 has been rolled out to 100% of engineers at Meta.
6. Balancing Reviewer Workload

### v2

@@ 3.1. Design and Feature Importance for RevRecV2
Prior works have mostly ranked reviewers based on a set of features and rules, e.g., (Asthana et al., 2019; Mirsaeedi and Rigby, 2020). Since reviewer recommendation is a ranking problem, we select learn-to-rank approach (Burges, 2010), LambdaMART. A pairwise comparison of each reviewer is made and this ordering is then used to determine the final order. In our case we use MART (Multiple Additive Regression Trees) as the function to determine the ordering of reviewers. The LambdaMART is the boosted tree version of LambdaRank, which is based on RankNet. All have proven to be very successful algorithms for solving real world ranking problems. An approximate description of how it works would take more space than this paper and is presented in (Burges, 2010). At Meta there is infrastructural support for LambdaMART models at scale using measures collecting data from internal databases on code review and coding.
Feature importance. In decision trees, feature importance is determined by how much each feature contributes to reducing the uncertainty in the target variable by the amount of reduction in the entropy that is achieved by splitting on a particular feature. Since LambdaMART uses boosted regression trees, the average reduction in entropy is calculated over the trees. The single-tree entropy is calculated by iterating over the non-leaf nodes of the tree and obtaining the weighted reduction in node entropy from the split at that node.
To train RevRecV2, we used the features shown in Table 1. The features were selected by the engineering team manually by using their domain knowledge of code review practices and their understanding of what metrics could be deployed in production. Specifically, they decided which features are most likely to contribute to the model’s prediction and then selected a subset that would would not slow down the recommender. Some features were included because the engineering team felt that they were inherently interesting. For example, although resigning from a diff has little predictive power, the engineers felt that a resignation demonstrated a clear example of a wrong reviewer suggestion. Since the number of features was small and prediction performance was stable, additional feature engineering was not conducted in this study especially since experiments showed substantial improvements over the baseline model. We trained the model on over 4.3 million instances of these features for reviewers. We evaluated the model on 1.1 million recommended candidates.
In Table 1, we show the feature importance across the 1.1 million diffs. The most important features that contribute to this accuracy improvement relate to the familiarity of the author with the review (time spent on the author’s prior diffs, 33.89%) and the reviewer participation rate (assigned reviews 31.28%). In contrast, code and review ownership are still useful but much less important. The number of files in the current diff that the candidate reviewer has authored in the past accounts for only 4.91% (authored diffs). The number of diffs that a reviewer has acted on and rejected have an importance of 10.10% (reviewer comments) and 11.56% (rejected diffs). The time spent when assigned a diff (explicit time spent) accounts for 6.69%. The total number of comments and number of resigned diffs have low importance with 1.57% and 0.00%, respectively. Our recommender clearly show that being a more active reviewer and having reviewed the author’s diff frequently in the past are much stronger predictors of who will actually perform a review than code ownership.
@@ 3.2. Design for RevRecWL
At Microsoft (Asthana et al., 2019), they found that initial feedback from developers indicated that reviewers from other projects complained that they were being incorrectly recommended. We made three compromises to convince the engineering team to run the experiment: (1) we set $\theta=0.1$ which reduces the influence of workload favoring expertise, (2) we only re-rank the top five suggestions from RevRecV2, and (3) instead of assigning the top ranked reviewer, we allow the author to select one of the top five reviewers. These latter two are important guarantees, first, people with low workload, but little experience will not be inadvertently added to the list eliminating the issue that Microsoft experienced. Second, the Top5 accuracy of RevRecWL = RevRecV2. Since the top five developers are displayed in Phabricator, this means in the most extreme case, the top developer from RevRecV2 will be displayed as the 5th developer, if they have high workload, but the same reviewers will be displayed only the order will change.
@@
(2) $\text{{RevRecWL}}(r)=\frac{\text{{RevRecV2}}(r)}{\text{Load}(r)}$
Instead of choosing a single measure of workload, we experiment with three types of workload weights, two are novel: the number of upcoming meetings in the next seven days, the reviewer activity in the last seven days (Asthana et al., 2019), and time spent reviewing in the last seven days. We calculate each over a seven day period because developer patterns on a particular day tend to vary, but the entire work week tends to be more predictable (Chen et al., 2022). In Section 6, we present our historical analysis to determine the best workload measure and our A/B test to determine if we can use RevRecWL in production.
@@ 3.3. Design for RecBystander
The bystander effect is a well known psychological problem that when a group of people are present, it is less likely that any individual will take action (Fischer et al., 2011). At Meta, developers work in a monorepo, so there are times when the author does not know which engineer to assign the review to. In other cases, the author might not mind and anyone on a particular team can do the review. We decided to randomly assign one of the top three reviewers from RevRecV2 to the review. The results for BystanderRecRnd are presented in Section 7.
4. Experimental Method and Outcome Metrics
Industry experiments are rarely reported in software engineering literature with most empirical studies being observational, where past data is analyzed via a regression analysis in an attempt to suggest causal relationships. In observational studies the treatment may be associated with some unobserved property of the subjects or tasks that affects the outcome i.e. allocation bias. Furthermore, the code recommender literature (e.g., (Thongtanunam et al., 2015)) assumes that the actual reviewer observed in the historic data is the gold standard (correct answer) for any proposed recommender. This ignores the possibility that the actual reviewer may not always represent the optimal choice and that the developer would not be affected by the recommendations they might be shown if a different recommender were to be used. This “backtesting” approach consists of calculating recommendations based on data predating the reviewer assignment and then checking if the actual reviewer is from among the TopN recommended reviewers.
A more reliable way to establish treatment effects is to run a randomized controlled experimental trial (Jadad and Enkin, 2007). In software engineering these trials are usually called A/B tests (see, for example, (Juristo and Moreno, 2013), for the discussion of various types of experimentation in software engineering). Per definitions in (Cook et al., 2002): an experiment is a study in which an intervention is deliberately introduced to observe its effects and in randomized experiments units are assigned to receive the treatment or an alternative condition by a random process. Industry experiments in software engineering are rare (Sjøberg et al., 2005; Juristo, 2016) for a number of reasons. For example, experiments involving creation of systems or features in parallel (see, e.g.,  (Anda et al., 2008)) are prohibitively expensive. Involving human subjects also requires addressing numerous ethical and privacy concerns. Some of the early experiments in software engineering did, however, involve code reviews (Porter et al., 1998).
Specifically, we are running experiments with a randomized intervention where each unit is in an A or B condition. The randomization assures that observed (and unobserved) context will be evenly distributed under the A or B conditions. In our case, each experiment includes an older recommender, condition A or baseline condition, and a new recommender, condition B. Like in clinical trials of drugs, for example, a drug should have an intended effect, such as reducing inflammation. Similarly, the new recommender was designed with a goal in mind, e.g. workload balancing. The drug, even if it is highly effective at reducing inflammation, may have undesired side effects, such as higher mortality and such side effects are recorded in clinical trials. If the side effects are stronger or more frequent in the experiment than in the baseline condition, the trial may need to be stopped. Similarly, we do not want to negatively impact our code review process, so we have guardrail metrics like how long a reviewer spends looking at a diff because we do not want people to start rubberstamping and spending less time on review because of our change in recommendations. If we see regressions in these guardrail metrics the experiment can be stopped early or will not be deployed to the entire company.
A/B test requires that a treatment (specific recommender) be assigned to subjects (engineers in parallel design or individual diffs in crossover design) at random with another option being the baseline (the existing recommender). The main advantage of random assignment is the removal of the allocation bias, i.e., it ensures that the distribution of the potential confounders are independent of the treatment. Also, the researcher can not assign treatments (and, thus, can not steer certain reviewers to certain treatments). Since the user interface in each treatment is exactly the same, the engineer is also not aware if they are in the baseline or treatment conditions so they can not intentionally act differently depending on the condition.
In contrast, if we were conducting an observational study, we would need to include numerous control variables, for example, the size and complexity of the diff (and of the underlying codebase), author and reviewer expertise, programming language, type of functionality being implemented, time of the year, day of the week, and numerous other factors that may affect the outcome and can (or in many contexts can not) be measured. Using randomization helps to control for both observed and unobserved variables. Results from A/B tests are, therefore, relatively easy to analyze. Because of the random assignment, on average, the only differences between recommenders are due to recommenders.
While software engineering experiments are rare, the software industry commonly uses the so-called A/B testing (Kohavi et al., 2020) to describe the evaluation protocol for, typically, user interface modifications. For example, when a user goes to a web page, they are randomly assigned either a regular version of a service or a modified version. Typically many different trials are ongoing at the same time so the randomization follows a factorial design to ensure that effects of each intervention (and, if possible, some of the interactions) can be estimated from the data. To implement A/B test for reviewer recommender, we ran a one-factor two-treatments controlled experiment as the two recommenders were the only distinction between the treatments. In many such experiments the user may see a visual difference and recognize that they are receiving experimental treatment and act differently, but in this case no such difference was observable, hence the treatment was “blinded” from the subject.
To conduct our experiments, we used an internal tool designed to support A/B testing. The tool is programmed to do randomization where, in some experiments, we used parallel design where each developer gets only one recommender, and in others we use a crossover design where each diff gets a randomly chosen recommender so that each developer gets assigned one recommender for some of the diffs and another for other diffs. The tool is then programmed to collect both goal and guardrail metrics and to run during a preset period of time. Once the experiment is complete, it presents an analysis page where average value and a confidence interval (at pval=0.01) are shown for each metric. The tool uses Welch’s t-test for unpaired samples to calculate the confidence interval. If appropriate permissions are secured, researchers may access raw collected data for usually 90 days but up to 6 months after the experiment. After that the data is deleted only the summaries (means, sample sizes, and standard deviations are available).
From the developers’ perspective, the review recommendation is a service that, when an engineer loads the diff page, gets automatically invoked and provides its recommendations. It is during that invocation that each diff may get a different recommender (or a developer keeps the same algorithm for subsequent diffs in case of parallel design). Once the page is loading, the developer may interact with the recommender (due to computational latency the recommender may not always appear immediately), or type in reviewer name directly. Hence, the record, in addition to the time of the action and the actual reviewer(s) chosen, also indicates if the recommender was interacted with (clicked on).
As in any experiments involving human subjects, the experimental design is reviewed to ensure the subjects will not be harmed and the tool used provides strong privacy protections. For example, the researcher does not have automatic access to the data that can be used to identify subjects, the data is provided by the tool in an aggregate form (means and standard deviations for each condition), and UID (user identifiable data) is completely removed after 90 days.
@@ 4.1. Goal and guardrail Outcomes Metrics
The vast majority of reviewer recommendation work does not release the recommender in production and only evaluates the accuracy on historical data. In contrast, the culture at Meta uses a sophisticated A/B test of every new feature and change. In Table 2 we show the recommenders that are being compared, the unit and size, and the time period of the experiment.
The team defines goal and guardrail metrics in advance of the experiment and then an A/B test is used to test the existing feature (the control or A condition) with the new feature (the test or B condition). The goal metrics measure the desired impact of the intervention. In contrast, guardrail metrics are the outcome metrics we do not want to negatively impact. In both cases, the null hypothesis or H0, is that nothing will change under the intervention. In Table 3, we introduce the metrics that we use as outcomes in this paper: TopN accuracy, TimeInReview, TimeSpent, Latency, Clicks on recommendations, and Workload. Depending on the experiment these outcomes will either be goal metrics or guardrails. For continuous data, we use a t-test, and for count data we use a Fisher test. Since workload is skewed we use a Wilcoxon signed-rank test. We often have multiple hypotheses, so we use a 99% confidence interval. We also report the actual p-value when the result is not statistically significant. We report the percentage point change rather than effect size measurements like small, medium, and large because engineers prefer percentage points as they can more easily discuss the magnitude and importance of the change in the context of the problem they are solving.
Table 2. We conduct three experiments. In each experiment we have a control recommender vs a test recommender. We report the experimental unit and when the experiment was run.
Experiment Description Expt. 1 Expt. 2 Expt. 3
Conditions: A vs B We conduct controlled experiments with a control or A condition vs a test or B condition RevRecV1 vs RevRecV2 RevRecV2 vs RevRecWL TeamOnly vs BystanderRecRnd
Experimental Unit The experimental unit is either the diff or the author. In either case, the unit is evenly divided across control and test conditions. 82k diffs 28k authors 12.5k authors
Experiment Date We cannot report the exact timeframe that we ran the experiment, but we report the season in which each experiment was conducted. Spring 2022 Winter 2023 Spring 2023
Table 3. For each experiment, we describe each outcome measure, indicate whether it is a Goal measure or a guardrail measure, and describe the statistical test that is appropriate for the type of data.
Outcome measures Description Expt. 1 Expt. 2 Expt. 3 Stat Test
TopN Accuracy How often does one of the recommended reviewers review the diff? We report Top1 and Top3 accuracy. Goal guardrail NA Fisher
TimeInReview The time when a diff is waiting for review or is being actively reviewed. It excludes the time when the author is reworking the code. We want to make sure that our recommender does not slow down reviewers. guardrail guardrail Goal t-test
TimeSpent The amount of time reviewers spend examining each diff. This is time spent on the review page for a diff. We want to make sure that reviewers are not rushing through reviews and rubberstamping diffs. guardrail guardrail guardrail t-test
Latency The amount of time from when the diff is published until the recommendations are displayed. We monitor the 90th and 99th percentile. The goal is to make sure that most diffs have recommendations produced by the time the diff page has loaded. Goal NA NA none
Clicks The number of times that an author selects one of the recommended reviewers. We want to know how often the recommendations are selected. Goal guardrail NA Fisher
Workload We measure the amount of time a candidate reviewer has spent looking at diffs in the last week. The distribution is skewed and our goal is to distribute review workload more evenly. NA Goal NA Wilcoxon signed-rank test
5. Expt. 1: Accuracy and Latency
@@
In Section 3, we examined the code reviewer recommendation literature. We use features from each category of measures we identified in the literature (Al-Zubaidi et al., 2020): code ownership, reviewer experience, reviewer participation rate, and familiarity with the author. We also use a novel measure of how long a reviewer has viewed the author’s changes in the last 90 days called TimeSpent.
@@
We conducted an A/B test experiment to understand if the new model is effective in production. Our experimental unit is the diff under review. Diffs in the control group, A, continued to receive reviewer recommendations from RevRecV1, the existing model in production. Diffs in the test group, B, received reviewer recommendations from our new model. Our experiment ran in late Spring of 2022, on 82k diff which were evenly and randomly split between the two conditions. We use a t-test on continuous data and a Fisher test on count data. Since we have multiple goal metrics we use a 99% confidence interval, i.e. $p<0.01$.
@@
Our goal metrics, the metrics we wanted to improve, were the Top1 and Top3 recommendation accuracy. TopN accuracy in this context is the fraction of times where at least one person belonging to the TopN predictions has either commented or acted, e.g., accepted, the diff. We also monitor when a recommendation is clicked, and we expect to see more accurate recommendations being clicked more often. Before we released our new recommender into production, we calculated the latency. In a dashboard we tracked the slowest diffs, i.e. the latency at the 90th and 99th percentiles. Because latency was measured outside of the experiment framework, we were not able to calculate statistical significance on this measure.
@@
Our guardrail metrics, the metrics we did not want to be negatively influenced, where the time diffs were in review (cycle time) and the TimeSpent that reviewers looked at each diff. We did not want to increase cycle time and we did not want recommended engineers to rubber stamp diffs and spend less time looking at them.
Experimental results: RevRecV1 vs RevRecV2
For Expt. 1, our goal was to improve the accuracy and latency. We had RevRecV1 as the control and RevRecV2 as the test. The unit was 82k diffs in the Spring of 2022. As we can see in Table 4, the experiment achieved our goals. We also found that the reduced latency improved the number of clicks on recommendations because they were displayed sooner. We did not want to see reviewers rushing reviews and did not expect to see a drop in how long reviews took.
Table 4. Expt. 1 the change in accuracy, clicks, latency, TimeInReview, and TimeSpent for RevRecV1 vs RevRecV2 on 82k diffs. Latency measures were calculated in a dashboard and no statistical test was conducted, the magnitude of the change is very large from an engineering perspective. Guardrail metrics are safety checks that we do not expect to see change.
Outcome Metric Type Expectation Change p-value
Top1 Accuracy Goal Increase $+9.14$ pp. $p\ll 0.01$
Top3 Accuracy Goal Increase $+14.19$ pp. $p\ll 0.01$
Clicks Goal Increase $+27.4\%$ $p\ll 0.01$
Latency p90 Goal Decrease $-14x$
Latency p99 Goal Decrease $-9x$
TimeInReview Guardrail No Increase $p=0.33$
TimeSpent Guardrail No Decrease $p=0.58$
Table 5. Improvement in Accuracy for RevRecV2. The improvement is statistically significant with $p\ll 0.01$.
Top-1 Top-3
RevRecV1 43.40 59.05
RevRecV2 52.54 73.24
Improvement 9.14 points 14.19 points
@@
Table 7. We report the number of diffs that the models were able to create recommendations for and how often the author clicked on the recommended reviewer. The improvement in clicked recommendations is statistically significant with $p\ll 0.01$.
@@
In Table 5, we see that the new model is much more accurate than RevRecV1. Looking at Top1 accuracy and Top3 accuracy we see 9.14 and 14.19 percentage point improvements, respectively, over RevRecV1. These results were statistically significant with $p\ll 0.01$.
When comparing latency of the new model to RevRecV1, we see major wins in Table LABEL:tabLatency. The 90th percentile latency decreased from 4.43s to 0.3s (a 14x improvement), and p99 latency decreased from 12.37s to 1.26s (a 9x improvement). This reduces latency largely because our new model uses measures that are stored in a database and does not require blame information to be calculated on the fly. As noted before, we tracked this information in a dashboard outside our experimental framework and were unable to calculate statistical significance. However, from a developer’s perspective the latency reduction of 14 times is substantial.
We also see that diffs in the test condition have more clicks of the suggested reviewer feature, with a 27.3% increase in the number of diffs where a suggested reviewer was added using the recommender reviewer component on the diff review page, see Table LABEL:tabUsage.
We did not see any statistically significant change in how long diffs were in review. Importantly, we did not see any statistically significant change in how long reviewers examined the code under review indicating that reviewers are not rubberstamping diffs. Based on these results, RevRecV2 has been rolled out to 100% of engineers at Meta since early Summer 2022.
In the A/B test, we saw a Top3 accuracy of 73%, a 14 point improvement over the previous model. We also saw a 14 times reduction in latency at p90. The recommendations were selected by authors 27% more often with the new model. These improvements were all statistically significant with $p\ll 0.01$. We did not see any statistically significant changes in the review cycle time or the reviewer viewing time. RevRecV2 was rolled out to 100% of engineers at Meta in early Summer 2022.
6. Expt. 2: Balancing Reviewer Workload
Engineers have different skillsets which will always skew reviewer workload (Rigby and Storey, 2011), it is still interesting to understand if a recommender that is aware of workload might distribute the load of reviewing and potentially drive down cycle time by suggesting the least busy expert. For example, Microsoft weighted their code reviewer recommender by a candidate’s workload and reduced the time it takes for a review (Asthana et al., 2019). A similar experiment at Ericsson (Strand et al., 2020) balanced reviewer workload, but was unsuccessful and was rolled back. As a result, we build upon the Microsoft work and we take the score from RevRecV2 that is now in production and weight it by the workload of a candidate as described in the design and methods Section 3.2.
Our goal metric is to reduce the median amount of time spent reviewing the reviewer who made an action on a diff. We expect a decrease in accuracy in Top1 accuracy. However, because we only re-rank reviewers, and still display all five reviewers, the same reviewers will be shown only if the order will have changed.
Our other guardrail metrics are the same as in the previous section. We do not want to see an increase in review cycle time. We want people to remain thorough in reviewing, so the time spent per diff should remain unchanged. We do not want to see fewer overall clicks on our recommendations.
6.1. Historical Backtest Results
To determine which type of workload is most effective at distributing reviews, we conduct a historical back test on 1.2 million historical diff reviews that were performed in early 2023. The backtest is followed by a production controlled experiment. We backtest three types of workload weights, the number of upcoming meetings in the next seven days, the reviewer activity in the last seven days (Asthana et al., 2019), and time spent reviewing in the last seven days. We calculate each over a seven day period because developer patterns on a particular day tend to vary, but the entire work week tends to be more predictable (Chen et al., 2022). We also measure the impact on the Top1 and Top3 accuracy.
The top portion of Table 8, shows the results. We see that the Top1 accuracy decreases dramatically using workload balancing by around -20 percentage points compared to RevRecV2. The workload differences are less consistent, with upcoming meetings performing poorly, and review activity and time spent performing well with respective decreases of -57% and -73%. The Top3 accuracy shows similar patterns with time spent having the largest drop in workload.
Since we are only re-ordering the reviewers suggested by RevRecV2 and are still displaying the Top5 recommended reviewers, we decided to use the most aggressive weight metric, time spent reviewing in the last seven days, because it had the largest drop in workload. We implemented the algorithm in production and gated it for an A/B test.